2. In a random sample of 26 people, the mean commute time to work was 32.2 minutes and the standard deviation was 7.1 minutes. Assume the population is normally distributed and use a t-distribution to construct a 80% confidence interval for the population mean "u". What is the margin of error of "u"? Interpret the results.
Answer #1

Given that, sample size ( n ) = 26

sample mean (ar x) = 32.2

sample standard deviation ( s ) = 7.1

confidence level = 80%

significance level (alpha) = 1 - 0.80 = 0.20

t-critical value significance level (alpha) = 0.20 with degrees of freedom = n - 1 = 26 - 1 = 25 is, a/2-1.316

We want to find, the margin of error ( E ),

7.1 E ta/1.3161.8324 Vn /26

Therefore, the margin of error of mu is 1.8324